OpenStudy (anonymous):

I f(x)= 3x-9 and g(x)= -2x+2 find ( f o g)(x)

OpenStudy (helder_edwin):

just follow the definition \[ (f\circ g)(x)=f(g(x)) \]

mathslover (mathslover):

Ok ! So first of all : http://www.regentsprep.org/Regents/math/algtrig/ATP7/compositionfunctions.htm Just to let u know about (f o g)x

OpenStudy (anonymous):

Still lost.

mathslover (mathslover):

Ok ! no problem : given : f(x) = 3x-9 and g(x) = 2-2x right ?